A boat traveled 210 miles downstream and back. The trip downstream took 10 hours. The trip back took 70 hours. What is the speed of the boat in still water? What is the speed of the current?
Answer by Alan3354(69443)

Here's a similar example.
--------------
A plane flew 720 miles in 3 hours with a tailwind. It would take 4 hours to travel the same distance with a headwind. Find the speed of the plane in still air and speed of wind.
-------------------------------
720/3 = 240 mph
720/4 = 180 mph
------------
The airspeed is the average (240 + 180)/2 = 210 mph
The windspeed is the difference = 30 mph